OpenStudy (anonymous):

-7x + 5y = -56 -21x – 2y = -168 Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.

OpenStudy (anonymous):

what do you need to do with it

OpenStudy (anonymous):

olve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.

OpenStudy (anonymous):

solve

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

okay do you have to show your work because sometimes they want you to solve it a certain way and i don't know what you have already learned

OpenStudy (anonymous):

no they just want you to show your work just in case you make a mistake

OpenStudy (anonymous):

i will let someone else try because I mess these up a lot there are many ways to do it

OpenStudy (anonymous):

ok thanks any ways

OpenStudy (kropot72):

-7x + 5y = -56...............(1) -21x - 2y = -168............(2) The elimination method requires us to get an equation that has only one variable. This equation can then be solved to find the value of the remaining variable. We can eliminate x by firstly multiplying both sides of equation (1) by 3 to give: -21x + 15y = -168............(3) If we subtract (3) from (2) we get -21x - (-21x) - 2y - 15y = -168 - (-168)...........(4) Simplifying (4) we get -17y = 0............(5) Now can you solve (5) to find the value of y?

OpenStudy (kropot72):

@artishaalston Are you there?